OpenStudy (anonymous):

Express the following in the form lgy=mX+c, where m and c are constants and X is a function of x. Determine the constants m and c and the function X. y=100x^1.5 okay i really can't get my head around this problem ugh

OpenStudy (paxpolaris):

\[\log(ab)=\log(a)+\log(b)\]\[\log(m^n)=n \cdot \log(m)\]

OpenStudy (paxpolaris):

\[\Large y=100x^{1.5}\] \[\Large \implies \log(y) =\log \left( 100x^{1.5} \right)\]

OpenStudy (paxpolaris):

then apply the rules

OpenStudy (anonymous):

ohhhhhh lgy=100+lgx^1.5 lgy=lg100+1.5lgx lgy=1.5X+2

OpenStudy (paxpolaris):

\[\checkmark\]